What regular polygon has an exterior angle measure of 18?
Answer by stanbon(57229)

What regular polygon has an exterior angle measure of 18?
---------------------------
The exterior angles always add up to 360 degrees.
Your problem:
(one ext. angle)*(number of sides) = 360
number of sides = 360/18 = 20
---------------------------------
The regular polygon has 20 sides.
==================================
Cheers,
Stan H.